OpenStudy (gabylovesyou):

Solve for x, given the equation Square root of x plus 9 − 4 = 1. x = 16, solution is not extraneous x = 16, solution is extraneous x = 34, solution is not extraneous x = 34, solution is extraneous

OpenStudy (gabylovesyou):

@SolomonZelman ik its x = 16 but idk if its extraneous or not

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle \sqrt{x+9}-4=1}\)

OpenStudy (solomonzelman):

yes, then x=16. extraneous solution is a solution that doesn't satisfy the original equation. When you plug in 16 for x , does that make the statement \(\large\color{slate}{\displaystyle \sqrt{x+9}-4=1}\) true?

OpenStudy (gabylovesyou):

yes

OpenStudy (gabylovesyou):

so its A

OpenStudy (solomonzelman):

that means it is not an extraneous solution.

OpenStudy (solomonzelman):

yes